# Non-Fermi liquid behaviour of CDW instabilities in fractionally-filled moiré flatbands

### v1

@@ III.1 One-loop Feynman diagrams
Figure 3: The one-loop diagrams for (a) the boson self-energy, (b) the fermion self-energy, (c) $\Psi\Psi$-vertex correction, and (d) $\bar{\Psi}\bar{\Psi}$-vertex correction. Curves with arrows represent the bare fermion propagator $G_{(0)}$, whereas the wiggly lines represent the dressed bosonic propagator $D_{(1)}$.
The value of $x_{g}$ tells us that the coupling constant $g$ attains a marginal character at the upper critical dimension $d_{c}=5/2$, which imples that $g$ is relevant for $d<5/2$ and irrelevant for $d>5/2$. Identification of the value of $d_{c}$ makes a controlled approximation possible because it allows us to describe the interacting phase perturbatively by using an $\epsilon$-expansion, where the actual physical dimension is expressed as $d=5/2-\epsilon$. Since the orginal theory lives in $d=2$ spatial dimensions, we need to set $\epsilon$-expansion at the end of the systematic perturbative expansion. In this subsection, we elucidate the expressions for the one-loop diagrams, which are shown in Fig. 3.
@@
The details of the intermediate steps are shown in Appendix B. We note that since the bare boson propagator $D_{(0)}(k)$ is independent of $\mathbf{K}$, the loop integrations involving them are ill-defined, unless one resums a series of diagrams that provides a nontrivial dispersion along these frequency and momentum components. Hence, in all loop calculations involving the boson propagators, we include the lowest-order finite correction from the one-loop self-energy. Therefore, we will use the dressed propagator
@@
which is equivalent to rearranging the perturbative loop-expansions such that the one-loop finite part of the boson self-energy, dependent on $\mathbf{K}$, is included at the zeroth order. We would like to emphasize that $\Pi_{1}(k)$ is the so-called Landau-damped term which leads to the signature $\text{sgn}(k_{0})\,|k_{0}|^{2/3}$-dependence of the fermion self-energy, characterizing the NFL behaviour in various quantum critical systems [23, 24, 25, 26, 27, 31, 32].
@@
Here, the logarithmic divergence of the self-energy (i.e., logathrithmically divergent in the Wilsonian cutoff $\Lambda$) has been parametrized by a gamma-function pole at $\epsilon=0$ (in the language of dimensional regularization).
Using steps similar to the evaluation of the bosonic and fermionic self-energies, it can be shown that the vertex corrections are zero.
@@ III.2 Stability of the non-Fermi liquid phase
We use the framework of the minimal subtraction scheme for extracting the renormalization group (RG) equations in terms of the beta-function of the relevant coupling constant. While the details have been relegated to Appendix D, here we show the final results leading to establising the fact that there emerges a stable fixed-point in the infrared (IR) at the QCP.
@@
using the notation of Appendix D. To this leading order correction, we find that $Z_{2}=Z_{3}$, and they do not get any correction from the loop integrals.
@@
respectively. We also define the beta function for the coupling constant $g$ as
@@
The sole purpose of the introduction of the ad hoc mass scale $\mu$ is to regularize the theory, thus eliminating the infinities emerging from the loop integrals of Feynman diagrams. However, since physical quantities must be independent of $\mu$, as $\mu$ is not really a parameter of the fundamental theory, the bare parameters must be independent of it as well. Imposing this condition, as well as the requirement that the regular (i.e., non-singular) parts of the final solutions are of the forms
@@
The above set of equations have been obtained by (1) demanding that $\frac{d}{d\ln\mu}\,(\mbox{bare quantity})=0$; (2) plugging in the values from Eqs. (III.2) and (III.2); (3) expanding in powers of $\epsilon$; and (4) matching the coefficients of the regular powers of $\epsilon$ in the resulting equations.
@@
$\frac{\beta_{g}}{g}=-\,\frac{\epsilon}{2}+\frac{\mathcal{U}_{1}\,\tilde{g}}{2}$ (25)
Our interest is in the behaviour at the IR energy scales and, hence, we determine the RG flows with respect to the logarithmic length scale $l$, which is given by the derivative
@@
The IR fixed-point is characterized by the value
$\tilde{g}^{*}\equiv\frac{\epsilon}{\mathcal{U}_{1}}\,.$ (27)
One can check that this is an IR-stable fixed-point [25, 26, 32]. At the fixed point, we obtain the dynamical critical exponent and the anomalous dimensions for the fermions and the bosons to be
$z^{*}=1+\frac{2\,\epsilon}{3}\,,\quad\eta_{\psi}^{*}=\eta_{\phi}^{*}=\frac{\epsilon}{2}\,$, (28)
@@ IV Summary and outlook
In this paper, we have investigated the nature of a putative quantum phase transition between an NFL and a putative CDW ordered phase. The origin of such a system lies in certain fractionally-filled bands of moiré superlattices. The behaviour is modelled by using bosonic fields, whose momenta are centred around finite wavevectors characteristic of the $\mathbf{M}$ points of a hexagonal mBZ. Thus, the bosons are capable of sourcing CDW instabilities by coupling with the fermions residing in the vicinity of the hot-spots. Especially, we have focussed at the QCP, where the bosons become massless, thus causing strong interactions which can bring about an NFL-behaviour. Analyzing the effecttive minimal action, we have found that a stable interacting fixed point results from the RG equations, analogous to the NFL systems in other systems capable of hosting CDW instabilities [21, 22].
A distinctive feature of an underlying triangular-lattice symmetry is that bosonic self-interactions in the form of cubic terms are allowed. This is unlike the case of a square lattice, where $\phi\rightarrow-\phi$ symmetry rules out such odd-powered terms. However, if we do retain the cubic terms, we end up with a bosonic action resembling the three-state Potts model in $(2+1)$ dimensions. In fact, the three-state Potts model by itself is not amenable to analytical solutions [in $(2+1)$ dimensions], and various properties remain inconclusive even with numerical solutions [33] (for example, the nature of the phase transition itself, or the value of the upper critical dimension). On top of it, after the bosons are coupled to fermions, it becomes impossible to obtain an analytical solution using the patch-theory formalism. The patch description, on the other hand, cannot be done away with, as it is necessary to capture the crucial behaviour of the fermions that, in the low-energy limit, they can scatter only tangentially to the FS at the hot-spots. Therefore, we have found the solutions neglecting the three-boson terms. In the future, one might attempt to work with the full theory, retaining the cubic bosonic self-interactions, using methods like functional renormalization group (fRG). Another avenue to explore is to study transport characteristics at the low-temperature regime, in the vicinity of the QCP, using the techniques developed in our earlier works involving NFLs [38, 39, 40, 41, 42]. This might be extended to the question of collective modes of the critical FSs, which are the well-defined contours hosting the hot-spots interacting with the critical bosons [43, 44, 45, 46].
@@ Appendix A Ordered states from exact diagonalization
Figure 4: (a) Many-body energy spectrum at filling $\nu=1/4$ for $N_{\text{s}}=32$ sites. The lowest 15 values are shown for each momentum sector and the energy is offset with respect to the lowest-enrgy state. (b) Pair-correlation function $G(\mathbf{r})$ in the finite-sized system showing the 2x2 CDW order. This figure is reproduced with permission from Raul Perea Causin [19].
While the analysis based on the fluctuating quantum metric already provides useful insights indicating the potential stability of CDW order, in the following we show the results from many-body exact diagonalization (ED) calculations [19], providing an idea tof the nature of the most stable phase. The focus is on a spin-valley polarized TLG system at a filling fraction equalling $\nu=1/4$, where CDW and Fermi-liquid phases are expected to compete, based on the quantum-metric arguments. Although there is the possibility of other competing phases arising in the topological band, such as fractional Chern insulators [16, 47, 11] and composite Fermi liquids [48, 49] (which can be obtained by switching the sign of the vertical electric field) at filling factors $\nu=1/3$ and $1/2$, and different magnetic orders (possible if the spin-valley polarization condition is relaxed [50]), those situations are not considered. The purpose of providing the unpublished results of the ED calculations is to serve as a benchmark for the QFT analysis that has been carried out in the main text.
@@
is diagonalized in a finite system of interacting electrons, where the semicolon indicates normal ordering and $\rho(\mathbf{q})$ is the density operator. Furthermore, $c_{\mathbf{k}}$ ($c^{\dagger}_{\mathbf{k}}$) denotes the electron annihilation (creation) operator acting on the band under consideration. Carrying out the numerics in a system with $N_{\text{s}}=32$ moiré sites, the many-body spectrum with $N_{\text{e}}=\nu\,N_{\text{s}}$ electrons (with $\nu$ set to $1/4$) is illustrated in Fig. 4(a). The result indeed shows a gapped fourfold-degenerate ground state consistent, featuring a CDW that breaks translation symmetry, enlarging the unit cell of the system by a factor of 4. In particular, the degenerate ground states are related to each other by momentum translations corresponding to the $\mathbf{M}$-points of the mBZ. The crystalline nature of the phase is further manifested in the structure factor,
@@
which displays prominent peaks at the $\mathbf{M}$-points of the mBZ, which demarcate the end points of the three $\mathbf{M}_{n}$ vectors, with $n\in\{1,2,3\}$. The pair-correlation function
@@
clearly shows the structure of the CDW oredering in the real-space lattice, as captured by Fig. 4(b). There, the unit cell corresponds to $2\times 2$ moiré cells. The identification of a CDW state as the natural instability of the $\nu=1/4$ phase (also dubbed as a generalized Wigner crystal in the context of moiré materials [7, 8]) thus serves a benchmark for our QFT analysis. However, we must caution the reader that, while the nature of some of the correlated phases can be accessed via ED, this method is limited to small systems — a large Fock space (e.g. due to valley and spin degrees of freedom or absence of particle-number conservation) presents a significant numerical challenge at reasonable system sizes.
@@ Appendix B One-loop boson self-energy
$\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ refer to the two options of contracting the internal fermionic lines. The above is obtained from the expansion of the interaction term to the second order, and hence, there is an extra overall factor of $2/2!=1$. The overall contribution is $\text{Tr}\left[\gamma_{0}^{T}\,G\,\gamma_{0}\,G^{T}\right]$. Thus, the one-loop boson self-energy [cf. Fig. 3(a)] is defined by
@@
The identities $\gamma_{d-1}^{T}=-\gamma_{0}\,\gamma_{d-1}\,\gamma_{0}$ and $\Gamma^{T}=-\gamma_{0}\,\Gamma\,\gamma_{0}$ lead to $\gamma_{0}\,G^{T}(k_{1})\,\gamma_{0}=-\,G^{T}(k_{1})$. Using the commutation relations between the gamma-matrices, we get
@@
The denominator of the last factor forces the dominant contribution to the integral to come from $u\sim-\,e_{q}/2$, in the regime $|{\mathbf{Q}}|\ll q_{d}^{2}$. Therefore, we focus on the regime where $e_{q}<0$, approximate $\sqrt{u}$ by $\sqrt{-\,e_{q}/2}$, and approximate the integral by extending the lower limit of the integral over $u$ to -$\infty$, and obtain
@@
We will evaluate the (d1d-1)-dimensional integral in $I_{1}(d,{\mathbf{Q}})$ by using the Feynman parametrization:
@@
The final expression, thus, turns out to be
@@ Appendix C One-loop fermion self-energy
The steps to compute these two parts have been explained in the next two subsections. Here, we state here the final result obtained on setting $d=d_{c}-\epsilon$, giving us the part $\propto 1/\epsilon$:
@@
In the above expressions, the logarithmic divergence of the self-energy (in the Wilsonian language) has been parametrized by a pole at $\epsilon=0$ (in the language of dimensional regularization).
@@ C.1 Computation of $\mathbf{\Gamma}$-dependent part
The leading order dependence of $\Sigma_{1}(k)$ on $\mathbf{K}$ can be extracted by setting the external momentum components $k_{d}$ and $k_{d-1}$ to zero. Hence, we will evaluate
@@
$=g^{2}\,\mu^{x_{g}}\int_{0}^{\infty}\frac{du}{\sqrt{u/2}}\int_{0}^{\infty}de_{q}\int_{-\infty}^{\infty}\frac{d^{d-1}\mathbf{Q}}{(2\,\pi)^{d+1}}\,\frac{\mathbf{\Gamma}\cdot(\mathbf{Q}-{\mathbf{K}})}{(\mathbf{Q}-{\mathbf{K}})^{2}+\left(u-e_{q}\right)^{2}}\,\frac{2}{2\,u+g^{2}\,\mu^{x_{g}}\,\beta_{d}\,|\mathbf{Q}|^{d-1}/\sqrt{e_{q}}}\,\,\left[\text{where }2\,u={q_{d}^{2}}\right],$ (48)
@@
The integral $I_{1}$ cannot be evaluated exactly and we need to make some reasonable approximations to extract the leading order corrections. We note that the first factor of the integrand tells us that the dominant contribution is concentrated around the region $|\mathbf{Q}|\sim|\mathbf{K}|$ and $u\sim e_{q}$. As for the second factor, the dominant contribution comes from $e_{q}\sim|\mathbf{Q}|^{2\,(d-1)/3}\sim|\mathbf{K}|^{2\,(d-1)/3}$. Since $|\mathbf{K}|^{2\,(d-1)/3}\gg|\mathbf{K}|$ for small $|\mathbf{K}|$ close to zero and 2(d1)/3<12\,(d-1)/3<1, we can substitute $u\sim e_{q}$ in the $\sqrt{u}$ factor in the overall denominator and the 2u2\,u term in the denominator of the second factor, and extend the lower limit of the integral over $u$ to -$\infty$, leading to
$I_{1}({\mathbf{K}})$ $\simeq g^{2}\,\mu^{x_{g}}\int_{-\infty}^{\infty}\frac{d^{d-1}\mathbf{Q}\,du}{(2\,\pi)^{d+1}}\int_{e_{q}>0}\frac{de_{q}}{\sqrt{e_{q}/2}}\,\frac{\mathbf{\Gamma}\cdot{\mathbf{Q}}}{\mathbf{Q}^{2}+u^{2}}\,\frac{2}{2\,e_{q}+g^{2}\,\mu^{x_{g}}\,\beta_{d}\,|\mathbf{Q}+{\mathbf{K}}|^{d-1}/\sqrt{e_{q}}}\quad\left[\text{shifting }u\rightarrow u+e_{q}\right]$
$=g^{2}\,\mu^{x_{g}}\int_{-\infty}^{\infty}\frac{d^{d-1}\mathbf{Q}\,du}{(2\,\pi)^{d+1}}\int_{e_{q}>0}de_{q}\,\frac{\mathbf{\Gamma}\cdot{\mathbf{Q}}}{\mathbf{Q}^{2}+u^{2}}\,\frac{2\,\sqrt{2}}{2\,e_{q}^{3/2}+g^{2}\,\mu^{x_{g}}\,\beta_{d}\,|\mathbf{Q}+{\mathbf{K}}|^{d-1}}$
@@
The integral blows up at $x_{g}=0\Rightarrow d=5/2$, which thus gives us the value of the upper critical dimension $d_{c}$. The fermion-boson coupling $e$ is irrelevant for $d>d_{c}$, relevant for $d<d_{c}$, and marginal for $d=d_{c}$. This allows us to access the strongly interacting state (NFL) perturbatively in a controlled approximation, using $d=5/2-\epsilon$, where $\epsilon$ serves as the perturbative/small parameter. In our dimensional regularization scheme, the divergence appears as $\sim\epsilon^{-1}$, with the $\Gamma\big(\frac{5-2\,d}{6}\big)$ factor having a pole at $d=d_{c}$. We also note that this term produces the behaviour of the fermion self-energy as $\sim\text{sgn}(k_{0})\,|k_{0}|^{2/3}$ at $d=2$, which matches with the uncontrolled RPA result [36, 51].

### v3

@@ III.1 One-loop Feynman diagrams
Figure 3: The one-loop diagrams for (a) the boson self-energy, (b) the fermion self-energy, (c) $\Psi\Psi$-vertex correction, and (d) $\bar{\Psi}\bar{\Psi}$-vertex correction. Curves with arrows represent the bare fermion propagator, $G$, whereas the wiggly lines in (b), (c), and (d) represent the dressed bosonic propagator, $D_{(1)}$. The wiggly line in (a) represents the bare bosonic propagator, $D_{(0)}$.
The value of $x_{g}$ tells us that the coupling constant $g$ attains a marginal character at the upper critical dimension $d_{c}=5/2$, which implies that $g$ is relevant for $d<5/2$ and irrelevant for $d>5/2$. Identification of the value of $d_{c}$ makes a controlled approximation possible because it allows us to describe the interacting phase perturbatively by using an $\epsilon$-expansion, where the actual physical dimension is expressed as $d=5/2-\epsilon$. Since the original theory lives in $d=2$ spatial dimensions, we need to set $\epsilon=1/2$ at the end of the systematic perturbative expansion. In this subsection, we elucidate the expressions for the one-loop diagrams, which are shown in Fig. 3.
@@
The details of the intermediate steps are shown in Appendix B. We note that since the bare boson propagator, $D_{(0)}(k)$, is independent of $\mathbf{K}$ and $k_{d-1}$. Thus, the loop integrations involving them are ill-defined, unless one resums a series of diagrams that provides a nontrivial dispersion along these frequency and momentum components. Hence, in all loop calculations involving the bosonic propagators, we include the lowest-order finite correction from the one-loop self-energy. Therefore, we will use the dressed propagator
@@
which is equivalent to rearranging the perturbative loop-expansions such that the one-loop finite part of the bosonic self-energy, dependent on $\mathbf{K}$, is included at the zeroth order. We would like to emphasize that $\Pi_{1}(k)$ is the so-called Landau-damped term which leads to the signature $\text{sgn}(k_{0})\,|k_{0}|^{2/3}$-dependence of the fermionic self-energy, characterising the NFL behaviour in various quantum critical systems [max-isn, max-sdw, Lee-Dalid, ips-uv-ir1, ips-uv-ir2, ips-fflo, ips-nfl-u1].
@@
Here, the logarithmic divergence of the self-energy (i.e., logarithmically divergent in the Wilsonian cutoff, $\Lambda$) has been parametrised by a gamma-function pole at $\epsilon=0$ (in the language of dimensional regularization).
Using steps similar to the evaluation of the bosonic and fermionic self-energies, it can be shown that the vertex corrections are non-divergent. Thus they do not contribute to the equations for renormalization.
@@ III.2 Stability of the non-Fermi liquid phase
For extracting the renormalization group (RG) equations in terms of the beta-function of the relevant coupling constant ($g$), we use the framework of the minimal subtraction scheme. While the details have been relegated to Appendix D, here we show the final results leading to establishing the fact that there emerges a stable fixed-point in the infrared (IR) at the QCP.
@@
using the notations of Appendix D. To this leading-order correction, we find that $Z_{2}=Z_{3}$, and they do not get any correction from the loop integrals.
@@
respectively. We also define the beta function for $g$ as
@@
The sole purpose of the introduction of the ad hoc mass scale, $\mu$, is to regularise the theory, thus eliminating the infinities emerging from the loop integrals of Feynman diagrams. However, since physical quantities must be independent of $\mu$, as $\mu$ is not really a parameter of the fundamental theory, the bare parameters must be independent of it as well. Imposing this condition, as well as the requirement that the regular (i.e., non-singular) parts of the final solutions are of the forms
@@
The above set of equations has been obtained by (1) demanding that $\frac{d}{d\ln\mu}\,(\mbox{bare quantity})=0$; (2) plugging in the values from Eqs. (III.2) and (III.2); (3) expanding in powers of $\epsilon$; and (4) matching the coefficients of the regular powers of $\epsilon$ in the resulting equations.
@@
$\frac{\beta_{g}}{g}=-\,\frac{\epsilon}{2}+\frac{\mathcal{U}_{1}\,\tilde{g}}{2}\,.$ (25)
Our interest is in the behaviour at the IR energy scales and, hence, we determine the RG flows with respect to the logarithmic length scale, $l$. This is captured by the differential equation,
@@
The IR fixed-point is characterised by the value
$\tilde{g}^{*}\equiv\frac{\epsilon}{\mathcal{U}_{1}}\,$, (27)
as a solution to $\frac{dg}{dl}=0$. One can check that this is an IR-stable fixed-point [Lee-Dalid, ips-uv-ir1, ips-nfl-u1]. At the fixed-point, we obtain the dynamical critical exponent and the anomalous dimensions for the fermions and the bosons to be
$z^{*}=1+\frac{2\,\epsilon}{3}\text{ and }\eta_{\psi}^{*}=\eta_{\phi}^{*}=\frac{\epsilon}{2}\,$, (28)
@@ IV Summary and outlook
In this paper, we have investigated the nature of a quantum phase transition between an FL and a putative CDW ordered phase. The origin of such a system lies in certain fractionally-filled bands of moiré superlattices. The behaviour is modelled by using bosonic fields, whose momenta are centred around finite wavevectors characteristic of the $\mathbf{M}$-points of a hexagonal mBZ. Thus, the bosons are capable of sourcing CDW instabilities by coupling with the fermions residing in the vicinity of the hot-spots. Especially, we have focussed at the QCP, where the bosons become massless, thus causing strong interactions which can bring about an NFL-behaviour. Analysing the effective minimal action, we have found that a stable interacting fixed-point results from the RG equations, analogous to the NFL systems in other systems capable of hosting CDW instabilities [ips-2kf, ips-cavity]. Although these systems arise in distinct contexts, their NFL-fixed-point behaviour turn out to be identical.
A distinctive feature of an underlying triangular-lattice symmetry is that bosonic self-interactions in the form of cubic terms are allowed. This is unlike the case of a square lattice, where $\phi\rightarrow-\phi$ symmetry rules out such odd-powered terms. However, if we do retain the cubic terms, we end up with a bosonic action resembling the three-state Potts model in $(2+1)$ dimensions. In fact, even the three-state Potts model by itself is not amenable to analytical solutions [in $(2+1)$ dimensions], and various properties remain inconclusive even with numerical solutions [wu-potts] (for example, the nature of the phase transition itself, or the value of the upper critical dimension). On top of it, after the bosons are coupled to fermions, it becomes impossible to obtain an analytical solution using the patch-theory formalism. The patch description, on the other hand, cannot be done away with, as it is necessary to capture the crucial behaviour of the fermions that, in the low-energy limit, they can scatter only tangentially to the FS at the hot-spots. Therefore, we have found the solutions neglecting the three-boson terms. In the future, one might attempt to work with the full theory, retaining the cubic bosonic self-interactions, using methods like functional renormalization group (fRG). Another avenue to explore is to study transport characteristics at the low-temperature regime, in the vicinity of the QCP, using the techniques developed in our earlier works involving NFLs [ips-subir, ips-hermann, ips-hermann2, ips-hermann3, ips-hermann-review]. This might be extended to the question of collective modes of the critical FSs, which are the well-defined contours hosting the hot-spots interacting with the critical bosons [else, ips-zero-mode, ips-kazi, ips-kazi2].
@@ Appendix A Ordered states from exact diagonalization
Figure 4: (a) Many-body energy spectrum at filling $\nu=1/4$ for $N_{\text{s}}=32$ sites. The lowest 15 values are shown for each momentum sector and the energy is offset with respect to the lowest-energy state. (b) Pair-correlation function, $G(\mathbf{r})$, in the finite-sized system showing a CDW order in real space. This figure is reproduced with permission from Raul Perea Causin [raul_unpub].
While the analysis based on the fluctuating quantum metric already provides useful insights indicating the potential stability ofa CDW order, in the following we show the results from many-body exact diagonalization (ED) [raul_unpub], providing an idea tof the nature of the most stable phase. The focus is on a spin- and valley-polarized TLG system at a filling-fraction equalling $\nu=1/4$, where CDW and FL phases are expected to compete, based on the quantum-metric arguments. Although there is the possibility of other competing phases arising in the topological band, such as fractional Chern insulators [abouelkomsan2020particle, repellin2020chern, xie2021fractional] and composite FLs [goldman2023zero, dong2023composite] (which can be obtained by switching the sign of the vertical electric field) at filling-factors $\nu=1/3$ and $1/2$, and different magnetic orders (possible if the spin- and valley-polarization conditions are relaxed [reddy2023fractional]), those situations are not considered. The purpose of providing the unpublished results of the ED calculations is to serve as a benchmark for the QFT analysis that has been carried out in the main text.
@@
is diagonalised in a finite system of interacting electrons, where the colons indicate normal-ordering of operators and $\rho(\mathbf{q})$ is the density operator. Furthermore, $c_{\mathbf{k}}$ ($c^{\dagger}_{\mathbf{k}}$) denotes the electron annihilation (creation) operator acting on the band under consideration. Carrying out the numerics in a system with $N_{\text{s}}=32$ moiré sites, the many-body spectrum with $N_{\text{e}}=\nu\,N_{\text{s}}$ electrons (with $\nu$ set to $1/4$) is illustrated in Fig. 4(a). The result indeed shows a gapped fourfold-degenerate ground state, featuring a CDW that breaks translation symmetry, enlarging the unit cell of the system by a factor of 4. In particular, the degenerate ground states are related to each other by momentum translations corresponding to the $\mathbf{M}$-points of the mBZ. The crystalline nature of the phase is further manifested in the structure factor,
@@
which displays prominent peaks at the $\mathbf{M}$-points of the mBZ. The $\mathbf{M}$-points demarcate the end points of the three $\mathbf{M}_{n}$ vectors, with $n\in\{1,2,3\}$. The pair-correlation function,
@@
clearly shows the structure of the CDW-ordering in the real-space lattice, as captured by Fig. 4(b). There, the unit cell corresponds to $2\times 2$ moiré cells. The identification of a CDW state as the natural instability of the $\nu=1/4$ phase (also dubbed as a generalised Wigner crystal in the context of moiré materials [regan2020mott, xu2020correlated]) thus serves a benchmark for our QFT analysis. However, we must caution the reader that, while the nature of some of the correlated phases can be accessed via ED, this method is limited to small systems — a large Fock space (e.g., due to valley and spin degrees of freedom, or absence of particle-number conservation) presents a significant numerical challenge at reasonable system-sizes.
@@ Appendix B One-loop boson self-energy
$\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ refer to the two options of contracting the internal fermionic lines. The above is obtained from the expansion of the interaction term to the second order, and hence, there is an extra overall factor of $2/2!=1$. The net contribution is thus $\text{Tr}\left[\gamma_{0}^{T}\,G\,\gamma_{0}\,G^{T}\right]$. Consequently, the one-loop boson self-energy [cf. Fig. 3(a)] is defined by
@@
The identities, $\gamma_{d-1}^{T}=-\gamma_{0}\,\gamma_{d-1}\,\gamma_{0}$ and $\Gamma^{T}=-\gamma_{0}\,\Gamma\,\gamma_{0}$, lead to $\gamma_{0}\,G^{T}(k_{1})\,\gamma_{0}=-\,G^{T}(k_{1})$. Using the commutation relations between the gamma-matrices, we get
@@
The denominator of the last factor forces the dominant contribution to the integral to come from $u\sim-\,e_{q}/2$, in the regime $|{\mathbf{Q}}|\ll q_{d}^{2}$. Therefore, we focus on the regime where $e_{q}<0$, approximate $\sqrt{u}$ by $\sqrt{-\,e_{q}/2}$, and approximate the integral by extending the lower limit of the integral over $u$ to -$\infty$. All these result in
@@
We will evaluate the (d1d-1)-dimensional integral in $I_{1}(d,{\mathbf{Q}})$ by using the Feynman parametrisation,
@@
The final expression turns out to be
@@ Appendix C One-loop fermion self-energy
The steps to compute these two parts have been explained in the next two subsections. Here, we state the final result obtained on setting $d=d_{c}-\epsilon$, giving us the part $\propto 1/\epsilon$:
@@
In the above expression, the logarithmic divergence of the self-energy (in the Wilsonian language) has been parametrised by a pole at $\epsilon=0$ (in the language of dimensional regularization).
@@ C.1 Computation of $\mathbf{\Gamma}$-dependent part
The leading-order dependence of $\Sigma_{1}(k)$ on $\mathbf{K}$ can be extracted by setting the momentum components, $k_{d}$ and $k_{d-1}$, to zero. Hence, we will evaluate
@@
$=g^{2}\,\mu^{x_{g}}\int_{0}^{\infty}\frac{du}{\sqrt{u/2}}\int_{0}^{\infty}de_{q}\int_{-\infty}^{\infty}\frac{d^{d-1}\mathbf{Q}}{(2\,\pi)^{d+1}}\,\frac{\mathbf{\Gamma}\cdot(\mathbf{Q}-{\mathbf{K}})}{(\mathbf{Q}-{\mathbf{K}})^{2}+\left(u-e_{q}\right)^{2}}\,\frac{2}{2\,u+g^{2}\,\mu^{x_{g}}\,\beta_{d}\,|\mathbf{Q}|^{d-1}/\sqrt{e_{q}}}\,\,\left(\text{where }2\,u={q_{d}^{2}}\right)$ (48)
@@
The integral $I_{1}$ cannot be evaluated exactly and we need to make some reasonable approximations to extract the leading-order correction. We note that the first factor of the integrand tells us that the dominant contribution is concentrated around the region $|\mathbf{Q}|\sim|\mathbf{K}|$ and $u\sim e_{q}$. As for the second factor, the dominant contribution comes from $e_{q}\sim|\mathbf{Q}|^{2\,(d-1)/3}\sim|\mathbf{K}|^{2\,(d-1)/3}$. Since $|\mathbf{K}|^{2\,(d-1)/3}\gg|\mathbf{K}|$ for small $|\mathbf{K}|$ close to zero and 2(d1)/3<12\,(d-1)/3<1, we can substitute $u\sim e_{q}$ in the $\sqrt{u}$ factor in the overall denominator and the 2u2\,u term in the denominator of the second factor, and extend the lower limit of the integral over $u$ to -$\infty$. This leads to
$I_{1}({\mathbf{K}})$ $\simeq g^{2}\,\mu^{x_{g}}\int_{-\infty}^{\infty}\frac{d^{d-1}\mathbf{Q}\,du}{(2\,\pi)^{d+1}}\int_{e_{q}>0}\frac{de_{q}}{\sqrt{e_{q}/2}}\,\frac{\mathbf{\Gamma}\cdot{\mathbf{Q}}}{\mathbf{Q}^{2}+u^{2}}\,\frac{2}{2\,e_{q}+g^{2}\,\mu^{x_{g}}\,\beta_{d}\,|\mathbf{Q}+{\mathbf{K}}|^{d-1}/\sqrt{e_{q}}}\quad\left(\text{shifting }u\rightarrow u+e_{q}\right)$
$=g^{2}\,\mu^{x_{g}}int_{-\infty}^{\infty}\frac{d^{d-1}\mathbf{Q}\,du}{(2\,\pi)^{d+1}}\int_{e_{q}>0}de_{q}\,\frac{\mathbf{\Gamma}\cdot{\mathbf{Q}}}{\mathbf{Q}^{2}+u^{2}}\,\frac{2\,\sqrt{2}}{2\,e_{q}^{3/2}+g^{2}\,\mu^{x_{g}}\,\beta_{d}\,|\mathbf{Q}+{\mathbf{K}}|^{d-1}}$
@@
Since $x_{g}=5/2-d$, $I_{1}({\mathbf{K}})$ blows up at $x_{g}=0\Rightarrow d=5/2$, which thus gives us the value of the upper critical dimension, $d_{c}$. The fermion-boson coupling ($g$) is irrelevant for $d>d_{c}$, relevant for $d<d_{c}$, and marginal for $d=d_{c}$. This allows us to access the strongly-interacting state (viz. NFL) perturbatively in a controlled approximation, using $d=5/2-\epsilon$, where $\epsilon$ serves as the perturbative/small parameter. In our dimensional regularization scheme, the divergence appears as $\sim\epsilon^{-1}$, with the $\Gamma\big(\frac{x_{g}}{3}\big)=\Gamma\big(\frac{5-2\,d}{6}\big)$ factor having a pole at $d=d_{c}$. We also note that this term produces the behaviour of the fermion self-energy as $\sim\text{sgn}(k_{0})\,|k_{0}|^{2/3}$ at $d=2$, which matches with the uncontrolled random-phase-approximation (RPA) result [metzner1, metzner2].